A ball is thrown vertically upward. The quantity which stays constant is a) displacement b)acceleration c)velocity d)speed e)momentum.
2) A ball is thrown from the ground into the air. At a height of 9.6m, the velocity is observed to be v= 8.0i + 6.5j in m/s. Find the maximum height the ball can reach.
1
Expert's answer
2017-07-06T10:28:07-0400
The acceleration stays constant or the answer b)

The second part:
Horizontal component of velocity is 8m/s
The velocity at 9.6m
v = sqrt(8^2 + 6.5^2)
v = 10.3 m/s
The whole energy at 9.6m is
E = 0.5mv^2 + mgh
E = m(0.5*10.3^2 + 9.8*9.6)
E = 147.125m
Energy at initial position is equal to the Energy at final position
147.125m = mgh + 0.5mv^2
Maximum height v = horizontal compnent of velocity
147.125m = m*9.8*H + 0.5*m*64
147.125 = 9.8H + 32
H = 11.8 m